OpenStudy (anonymous):

Solve the equation. 78 = -2(m + 3) + m

OpenStudy (anonymous):

please explain..

OpenStudy (compassionate):

Use PEMDAS (Parenthesis, Exponents, multiplication, division, addition, subtraction) We have parenthesis, but (m + 3) cannot be solved because they are not liket erms We have no exponents There is some multiplication we can do. We can distribute 2 78 = -2(m + 3) + m 78 = -2m - 3 + m Now we move onto addition/subtraction. We see we can add -2m and m because they are like terms 78 = -1m - 3 Now, we need to perform the opposite operation to isolate -m. That is, by adding 3 to both sides 78 + 3 = -1m 81 = -1m Now we simply divide by -1. So 81/-1 = ??

OpenStudy (anonymous):

A. – 28 B. – 42 C. – 72 D. – 84

OpenStudy (wolf1728):

78 = -2(m + 3) + m 78 = -2m -6 +m 78 = -m -6 84 = -m m = -84

OpenStudy (anonymous):

@wolf1728 i did it like you, but how did you get the 84?

OpenStudy (anonymous):

oh you i see , you added 6. nevermind. lol thankyouuu

OpenStudy (wolf1728):

Adding +6 to both sides - oh you see? okay

OpenStudy (anonymous):

thanks

OpenStudy (wolf1728):

u r welcome